40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Mon Jul 26 01:30:42 2021, Koji, Summary, Cryo vacuum chamber, Prep for the 2nd cooling of the suspension 
    Reply  Mon Jul 26 22:49:00 2021, Koji, Summary, Cryo vacuum chamber, Aquadag painting 10x
       Reply  Wed Jul 28 00:59:47 2021, Koji, Summary, Cryo vacuum chamber, The test mass successfully suspended P_20210727_154143.jpegP_20210727_190356.jpegP_20210727_190426.jpegP_20210727_190543.jpeg
          Reply  Thu Jul 29 00:42:38 2021, Koji, Summary, Cryo vacuum chamber, The test mass successfully suspended 
             Reply  Thu Jul 29 13:11:17 2021, Koji, Summary, Cryo vacuum chamber, Cooling progress: Update temp_log_cool_down_20210728_1830.pdfcooling_model1.pdfcooling_model2.pdfOSEM_cooling.pdf
                Reply  Fri Jul 30 12:22:56 2021, Koji, Summary, Cryo vacuum chamber, Cooling curve comparisons comparison_cold_head.pdfcomparison_inner_shield.pdfcomparison_test_mass.pdf
                   Reply  Sun Aug 1 22:22:00 2021, Koji, Summary, Cryo vacuum chamber, Cooling update temp_log_cool_down_20210728_1830.pdfcooling_meas.pdfOSEM_cooling.pdfcooldown_210728.zip
                      Reply  Mon Aug 2 13:25:32 2021, Stephen, DailyProgress, Cryo vacuum chamber, Warmup started 02 August 
                         Reply  Tue Aug 3 23:56:00 2021, Koji, DailyProgress, Cryo vacuum chamber, Warmup started 02 August P_20210804_000247.jpgP_20210803_235421.jpg
Message ID: 2633     Entry time: Tue Aug 3 23:56:00 2021     In reply to: 2630
Author: Koji 
Type: DailyProgress 
Category: Cryo vacuum chamber 
Subject: Warmup started 02 August 

- Confirmed the heating stopped in the evening -> The heater was deactivated @~23:00

- Made some measurements and checks - the oplev spot was approximately on the center of the QPD before warming up. Now it is ~4mm above the center (note that the QPD size is 0.5" in dia) (Attachment 2). This corresponds to ~2mrad misalignment.

- Dismantled the OSEM electronics and power supply from the table. The electronics were salvaged into the OMC lab -> to be returned to the 40m.

- A 2" Al mirror package was brought to the OPLEV periscope so that the gold mirror (too thin) can be replaced. (Attachment 1)

Attachment 1: P_20210804_000247.jpg  773 kB  Uploaded Wed Aug 4 02:09:03 2021  | Hide | Hide all
P_20210804_000247.jpg
Attachment 2: P_20210803_235421.jpg  3.150 MB  Uploaded Wed Aug 4 02:12:05 2021  | Hide | Hide all
P_20210803_235421.jpg
ELOG V3.1.3-